# Continuous-Time Value Iteration for Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Existing reinforcement learning (RL) methods struggle with complex dynamical systems that demand interactions at high frequencies or irregular time intervals. Continuous-time RL (CTRL) has emerged as a promising alternative by replacing discrete-time Bellman recursion with differential value functions defined as viscosity solutions of the Hamilton–Jacobi–Bellman (HJB) equation. While CTRL has shown promise, its applications have been largely limited to the single-agent domain. This limitation stems from two key challenges: (i) conventional solution methods for HJB equations suffer from the curse of dimensionality (CoD), making them intractable in high-dimensional systems; and (ii) even with HJB-based learning approaches, accurately approximating centralized value functions in multi-agent settings remains difficult, which in turn destabilizes policy training. In this paper, we propose a CT-MARL framework that uses physics-informed neural networks (PINNs) to approximate HJB-based value functions at scale. To ensure the value is consistent with its differential structure, we align value learning with value-gradient learning by introducing a Value Gradient Iteration (VGI) module that iteratively refines $\nabla_{x}V$ along trajectories. This improves gradient fidelity, in turn yielding more accurate values and stronger policy learning. We evaluate our method using continuous-time variants of standard benchmarks, including multi-agent particle environment (MPE) and multi-agent MuJoCo. Our results demonstrate that our approach consistently outperforms existing continuous-time RL baselines and scales to complex multi-agent dynamics.
@@ 1 Introduction
RL has achieved remarkable success in a range of single- and multi-agent interaction tasks, including robotic manipulation (Brunke et al., 2022), strategy games (Vinyals et al., 2019), wireless communications (Feriani and Hossain, 2021; Wang et al., 2023) , and traffic coordination (Haydari and Yılmaz, 2020). Most existing RL methods model these interactions in discrete time, where Bellman backup is computed at a fixed time interval. However, discrete-time RL (DTRL) is not well-suited for real-world scenarios that often demand high-frequency decision-making or operate at arbitrary, non-uniform time intervals (e.g., financial trading (Shavandi and Khedmati, 2022) and autonomous driving (Kiran et al., 2021)). Specifically, DTRL methods tend to generalize poorly when deployed under time resolutions that differ from training, leading to suboptimal control and stability issues (Wang et al., 2020; Shilova et al., 2024). To address these limitations, CTRL has emerged as an alternative to learn value functions in continuous time (Doya, 2000b; Rubanova et al., 2019). However, existing works only focus on single-agent settings and have not yet been widely explored for multi-agent scenarios. In multi-agent domains, each agent must not only interact with the environment but also coordinate with other agents, all while coping with the chllenges like non-stationarity introduced by simultaneously learning policies. Within this context, our paper investigates a novel approach to solving CT-MARL problems. To better understand the limitations of DTRL and advantages of CTRL for multi-agent scenarios, we present a didactic case as shown in Fig. 1. In this simple continuous-time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2. In contrast, our CT-MARL algorithm closely follows the ground-truth trajectory, maintains high returns, and generates accurate control actions for both agents.
@@
Unlike Bellman operator-based DTRL (Bellman, 1966), CTRL leverages HJB PDEs to compute differential value functions (Mukherjee and Liu, 2023; Shilova et al., 2024). However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or discretization) suffers from CoD in high-dimensional dynamical systems (Bellman et al., 1965), especially where the computational complexity grows exponentially with the state dimension in multi-agent systems. PINNs have emerged as a powerful tool to circumvent CoD (Weinan et al., 2021), and offer convergence guarantees for problems with smooth solutions (Shin et al., 2020; Ito et al., 2021). To approximate the solutions of HJB PDEs, PINNs translate the underlying physics law (e.g., PDEs) along with boundary conditions into the loss functions to refine networks.
However, even with PINNs to alleviate the CoD in CT-MARL, learning accurate centralized value functions under the centralized training decentralized execution (CTDE) framework remains challenging. The inherent non-stationarity across agents makes value learning unstable (Yu et al., 2022), and standard PINNs, which rely solely on PDE and boundary residuals, often yield biased or noisy value gradients (Zhang et al., 2024b). These inaccuracies propagate into the value function, undermining policy quality and stability. To address this limitation, we propose a novel learning approach that combines PINN and VGI information, which is used to optimize the value learning. The PINN component ensures the value approximations satisfy the HJB PDEs, while VGI iteratively propagates and refines the value gradient approximations along the sampled trajectories. This integration yields significantly more accurate value-gradient information, which enhances value approximation accuracy and ultimately improves policy learning.
Our work makes the following contributions. (1) We leverage PINNs to approximate differential value functions and apply them to solve continuous-time multi-agent reinforcement learning problems, which have rarely been explored by previous studies. (2) We introduce a novel value-gradient iteration term that dynamically refines the approximations of value gradients during training. This setting improves the computational accuracy of the value gradients, accelerates learning convergence, and leads to highly accurate value approximations, enabling efficient policy learning. (3) We create continuous-time versions of two standard MARL benchmarks, the continuous-time MPE and continuous-time multi-agent MuJoCo. The results demonstrate that our method consistently surpasses other current CTRL baselines, highlighting the advantages of precise value-gradient learning in high-dimensional multi-agent systems.
@@ 2.1 Continuous-Time Reinforcement Learning
CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings Jia and Zhou (2022); Lee and Sutton (2021); Wallace and Si (2023); Yang et al. (2021); Wiltzer et al. (2022); Vamvoudakis and Lewis (2010); Bian and Jiang (2021); Faradonbeh and Faradonbeh (2023). These studies aim to optimize policies in continuous-time domains, which avoids time discretization and achieves higher fidelity in control applications such as robotics and navigation. For instance, Bian and Jiang (2021) proposes a continuous-time value iteration algorithm to solve the HJB equation without the need for a stabilizing initial policy. Similarly, Faradonbeh and Faradonbeh (2023) designs an actor-critic network for solving the infinite horizon HJB equations with theoretical guarantees. Lee and Sutton (2021) designs two policy iteration algorithms that compute comprehensive solutions to HJB equations. In contrast, Jia and Zhou (2022) introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization. Yang et al. (2021) develops a robust actor–critic framework for continuous-time nonlinear systems using unmodeled dynamics. Shilova et al. (2024) utilizes PINNs with $\epsilon$-scheduling iteration to approximate the value function and empirically outperforms SOTA RL methods in continuous-time settings. Yildiz et al. (2021) integrates neural ordinary differential equations (ODEs) to model uncertainty in state evolution using Bayesian inference, and further introduces a continuous-time actor-critic algorithm that circumvents challenges such as Q-function vanishment and poor discretization. CT-MARL remains relatively underexplored compared to the substantial studies for single-agent settings. Luviano and Yu (2017) solves multi-agent pathfinding problem using fuzzy Q-learning, while Jiang et al. (2023) proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems. Beyond these examples, only a limited number of studies have addressed CT-MARL, which highlights the importance of our contributions.
@@ 2.2 Solving HJB Equations via PINNs
In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations Crandall and Lions (1983), which are the first-order nonlinear parabolic PDEs. However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD Osher et al. (2004); Osher and Shu (1991). Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth Weinan et al. (2021). PINNs often approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residual Han and Long (2020); Han et al. (2018), PDE residual Bansal and Tomlin (2021); Zhang et al. (2024b), and supervised data derived from numerical solvers Nakamura-Zimmerer et al. (2021); Zhang et al. (2024a). Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments Mukherjee and Liu (2023). However, solving CT-MARL problems through the integration of PINNs and RL remains an open and unexplored area of research.
@@ 3.1 Problem Formulation
In this paper, we focus on multi-agent cooperative settings. Following the continuous-time control system framework Doya (2000a); Yildiz et al. (2021); Lee and Sutton (2021), we formulate the continuous-time multi-agent problem as a tuple:
@@ 3.2 HJB and Policy Learning
First, we present the following Lemma to show that the optimal value function $V(x,t)$, as defined in Eq. (2), is the optimal solution to HJB PDEs Crandall and Lions (1983) in the context of cooperative multi-agent settings.
@@
To approximate differential value functions, we aim to solve the HJ PDEs given in Eq. (3). However, solving such PDEs through numerical methods becomes computationally intractable for problems with state dimensions higher than 6 Bui et al. (2022). To address this challenge, we adopt PINNs, which approximate value functions to satisfy both PDE residuals and boundary conditions. Specifically, we define HJB PDE residual as
@@ 3.2.2 Policy Learning
Assume the one-step Q-function over a short interval $\delta t>0$ be
@@ 3.3 Value Gradient Iteration Module
The performance of continuous-time control policies depends critically on the accuracy of the value, which in turn depends not only on the precision of its own approximation but also on the correctness of the value gradient $\nabla_{x}V(x,t)$. Recent studies have demonstrated that the accuracy of the value directly affects the computational closed-form control inputs Zhang et al. (2024b); Mukherjee and Liu (2023); Eberhard et al. (2024). However, the gradients obtained from standard PINN training are often noisy or misaligned with true trajectory behavior. Although prior work Zhang et al. (2024b) improves value gradient approximations using supervised data, such ground-truth gradients are often inaccessible or prohibitively expensive to compute in complex, high-dimensional multi-agent settings. To address this, we introduce the VGI module that iteratively refines the value gradient approximations during training.
@@ 3.4.1 Dynamics Model and Reward Model
In a continuous-time setting, the true dynamics are given by $\dot{x}=f(x,u)$, but directly learning $f$ via $\frac{x_{t+\Delta t}-x_{t}}{\Delta t}$ as a supervision target is pretty unstable in practice. Instead, we adopt a discrete-time model-based approach Sutton (1991); Kaiser et al. (2019); Hafner et al. (2019) that we train a neural network $f_{\psi}(x_{t},u_{t},\Delta t)$ to predict the next state $x_{t+\Delta t}$ via
@@
After learning $f_{\psi}$, we recover the continuous-time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t},\Delta t)-x_{t}}{\Delta t}$.
@@ 3.4.2 Anchor Loss for Critic Network
Here $\lambda_{\text{term}},\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE fidelity, terminal supervision, value-bootstrap anchoring, and gradient consistency. In practice, we jointly train the reward model, dynamics model, and value network using the data from the current trajectory. The detailed training process is listed in Appendix A.6.
@@ 4 Experimental Results
We evaluate our Value Iteration via PINN (VIP) method on two continuous-time multi-agent benchmarks: MPE Lowe et al. (2017b) and multi-agent MuJoCo Todorov et al. (2012); Peng et al. (2021). In addition, we design a didactic benchmark, coupled oscillator (see details in Appendix B.1), to analyze value gradient approximations. This case study is easy to follow and enables the numerical computation of true values and their gradients, which provides a clear and interpretable setting to validate the effectiveness of VIP. Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperforms existing continuous-time RL baselines in these environments? (2) VGI ablation: How much does the VGI module contribute to final performance and training stability? (3) PINN design choice: How does activation function choice and loss term weighting in critic network affect the performance of VIP? (4) Time discretization impact: How well do discrete-time and continuous-time methods perform under arbitrary or unfixed time intervals?
@@ 4.1 Baseline Methods
To evaluate our CT-MARL framework VIP, we compare against three continuous-time policy iteration baselines and include an ablated variant of our method without VGI: CT-MBRL (ODE) Yildiz et al. (2021): A continuous-time model-based RL approach that learns system dynamics via Bayesian neural ODEs. This method uses an actor-critic framework to approximate state-value functions and derive the continuous-time optimal policies.
@@
Differential Policy Iteration (DPI) Lee and Sutton (2021): A model-based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations. Integral Policy Iteration (IPI) Lee and Sutton (2021): A partially model-free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement. In our experimental settings, we discretize the integral, roll out trajectories to accumulate rewards, and fit a policy to minimize the resultant value functions. Ablation (w/o VGI): An ablated version of VIP without VGI, which isolates the efficacy of value gradient refinement.
@@ 4.2 Results Analysis
Model Performance. We evaluate all RL methods in MPE and MuJoCo environments using five random seeds and report the mean cumulative reward curves in Fig. 2 and 4. The results show that VIP with VGI consistently converges fastest and achieves the highest final return across all tasks, which empirically validates the efficacy of integrating PINNs with RL. As HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations Krishnapriyan et al. (2021); Zhang et al. (2024b).
@@
To address this limitation, we incorporate the anchor and VGI loss terms in Eq. (18), which capture the landscape of value and its gradient so that PINNs converge to the true values. Ablation results further confirm the importance of VGI: removing VGI leads to significantly lower cumulative rewards across all experiments. This observation is consistent with the conclusion in the previous studies Zhang et al. (2024a), which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement.
@@
Choice of PINN Design. We first examine the impact of activation function choice when incorporating PINN-based losses into critic network refinement. Specifically, we train VIP using ReLU and Tanh activations in both MuJoCo and MPE environments and report the accumulated rewards in Fig. 5. The results show that the VIP with Tanh consistently achieves higher accumulated rewards than the one with ReLU across all tasks. This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies Bansal and Tomlin (2021); Zhang et al. (2024b). The good performance of Tanh can be attributed to its smoothness and differentiability, which are particularly important when using PINNs to solve PDEs. PINNs typically use fully-connected network architectures and rely on auto-differentiation to compute value gradients for PDEs. The PINN residuals, including PDEs and boundary conditions, are further optimized using gradient descent. Smooth activation functions like Tanh support stable and accurate gradient flow throughout training, enabling more effective value approximations.
@@
We also investigate the impact of weight parameters for PINN loss terms during VIP training. Specifically, we evaluate three configurations in Eq. (18): 1) balanced weights for all loss terms; 2) a large weight for boundary condition while keeping the others balanced; 3) a large weight for HJB residual with balanced weights for the remaining terms. As shown in Fig. 7, the best accumulated reward performance is achieved when all loss terms are properly balanced. Imbalanced weight settings yield stiffness in the training dynamics of PINNs Wang et al. (2021), which makes the dominant loss term (with the largest weights) converge faster than the others Wang et al. (2022).
@@
In our experiments, such an imbalance causes VIP only to satisfy the PDEs or boundary conditions during training, ultimately leading to poor value approximations. Lastly, we evaluate the robustness of VIP and MADDPG Lowe et al. (2017a) by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts. Fig. 8 illustrates that VIP maintains a nearly constant return across different time intervals, whereas MADDPG’s performance degrades significantly as the interval increases. This result highlights the advantage of VIP in continuous-time multi-agent scenarios.

### v2

@@ Abstract
Existing reinforcement learning (RL) methods struggle with complex dynamical systems that demand interactions at high frequencies or irregular time intervals. Continuous-time RL (CTRL) has emerged as a promising alternative by replacing discrete-time Bellman recursion with differential value functions defined as viscosity solutions of the Hamilton–Jacobi–Bellman (HJB) equation. While CTRL has shown promise, its applications have been largely limited to the single-agent domain. This limitation stems from two key challenges: (i) conventional solution methods for HJB equations suffer from the curse of dimensionality (CoD), making them intractable in high-dimensional systems; and (ii) even with HJB-based learning approaches, accurately approximating centralized value functions in multi-agent settings remains difficult, which in turn destabilizes policy training. In this paper, we propose a CT-MARL framework that uses physics-informed neural networks (PINNs) to approximate HJB-based value functions at scale. To ensure the value is consistent with its differential structure, we align value learning with value-gradient learning by introducing a Value Gradient Iteration (VGI) module that iteratively refines $\nabla_{x}V$ along trajectories. This improves gradient fidelity, in turn yielding more accurate values and stronger policy learning. We evaluate our method using continuous‑time variants of standard benchmarks, including multi‑agent particle environment (MPE) and multi‑agent MuJoCo. Our results demonstrate that our approach consistently outperforms existing continuous‑time RL baselines and scales to complex multi-agent dynamics.
@@ 1 Introduction
RL has achieved remarkable success in a range of single- and multi-agent interaction tasks, including robotic manipulation [5], strategy games [39], wireless communications [12, 44] , and traffic coordination [16]. Most existing RL methods model these interactions in discrete time, where Bellman backup is computed at a fixed time interval. However, discrete‑time RL (DTRL) is not well-suited for real-world scenarios that often demand high-frequency decision-making or operate at arbitrary, non-uniform time intervals (e.g., financial trading [33] and autonomous driving [21]). Specifically, DTRL methods tend to generalize poorly when deployed under time resolutions that differ from training, leading to suboptimal control and stability issues [41, 34]. To address these limitations, CTRL has emerged as an alternative to learn value functions in continuous time [9, 32]. However, existing works only focus on single‑agent settings and have not yet been widely explored for multi-agent scenarios. In multi-agent domains, each agent must not only interact with the environment but also coordinate with other agents, all while coping with the chllenges like non-stationarity introduced by simultaneously learning policies. Within this context, our paper investigates a novel approach to solving CT-MARL problems. To better understand the limitations of DTRL and advantages of CTRL for multi-agent scenarios, we present a didactic case as shown in Fig. 1. In this simple continuous‑time control task, DT-MARL fails to accurately approximate the true value functions, leading to incorrect control actions, particularly for agent 2. In contrast, our CT‑MARL algorithm closely follows the ground‑truth trajectory, maintains high returns, and generates accurate control actions for both agents.
@@
Unlike Bellman operator-based DTRL [2], CTRL leverages HJB PDEs to compute differential value functions [27, 34]. However, solving HJB PDEs through conventional approaches (e.g., dynamic programming or discretization) suffers from CoD in high-dimensional dynamical systems [3], especially where the computational complexity grows exponentially with the state dimension in multi-agent systems. PINNs have emerged as a powerful tool to circumvent CoD [45], and offer convergence guarantees for problems with smooth solutions [35, 17]. To approximate the solutions of HJB PDEs, PINNs translate the underlying physics law (e.g., PDEs) along with boundary conditions into the loss functions to refine networks.
However, even with PINNs to alleviate the CoD in CT-MARL, learning accurate centralized value functions under the centralized training decentralized execution (CTDE) framework remains challenging. The inherent non-stationarity across agents makes value learning unstable [49], and standard PINNs, which rely solely on PDE and boundary residuals, often yield biased or noisy value gradients [51]. These inaccuracies propagate into the value function, undermining policy quality and stability. To address this limitation, we propose a novel learning approach that combines PINN and VGI information, which is used to optimize the value learning. The PINN component ensures the value approximations satisfy the HJB PDEs, while VGI iteratively propagates and refines the value gradient approximations along the sampled trajectories. This integration yields significantly more accurate value‑gradient information, which enhances value approximation accuracy and ultimately improves policy learning.
Our work makes the following contributions. (1) We leverage PINNs to approximate differential value functions and apply them to solve continuous‑time multi‑agent reinforcement learning problems, which have rarely been explored by previous studies. (2) We introduce a novel value‑gradient iteration term that dynamically refines the approximations of value gradients during training. This setting improves the computational accuracy of the value gradients, accelerates learning convergence, and leads to highly accurate value approximations, enabling efficient policy learning. (3) We create continuous‑time versions of two standard MARL benchmarks, the continuous-time MPE and continuous-time multi-agent MuJoCo. The results demonstrate that our method consistently surpasses other current CTRL baselines, highlighting the advantages of precise value‑gradient learning in high‑dimensional multi‑agent systems.
@@ 2.1 Continuous-Time Reinforcement Learning
CTRL has received increasing attention in recent years, however, most existing works focus on the single-agent settings [18, 23, 40, 47, 46, 38, 4, 11]. These studies aim to optimize policies in continuous-time domains, which avoids time discretization and achieves higher fidelity in control applications such as robotics and navigation. For instance, [4] proposes a continuous-time value iteration algorithm to solve the HJB equation without the need for a stabilizing initial policy. Similarly, [11] designs an actor-critic network for solving the infinite horizon HJB equations with theoretical guarantees. [23] designs two policy iteration algorithms that compute comprehensive solutions to HJB equations. In contrast, [18] introduces a temporal–difference learning based algorithm to deal with continuous-time problems via discretization. [47] develops a robust actor–critic framework for continuous-time nonlinear systems using unmodeled dynamics. [34] utilizes PINNs with $\epsilon$-scheduling iteration to approximate the value function and empirically outperforms SOTA RL methods in continuous-time settings. [48] integrates neural ordinary differential equations (ODEs) to model uncertainty in state evolution using Bayesian inference, and further introduces a continuous-time actor-critic algorithm that circumvents challenges such as Q-function vanishment and poor discretization. CT-MARL remains relatively underexplored compared to the substantial studies for single-agent settings. [26] solves multi-agent pathfinding problem using fuzzy Q-learning, while [19] proposes a model-based value iteration algorithm tailored for continuous-time multi-agent systems. Beyond these examples, only a limited number of studies have addressed CT-MARL, which highlights the importance of our contributions.
@@ 2.2 Solving HJB Equations via PINNs
In single-agent optimal control or multi-agent cooperative settings, value functions are characterized as the viscosity solutions to HJB equations [7], which are the first-order nonlinear parabolic PDEs. However, solving HJB equations with conventional numerical methods is computationally intractable in high-dimensional settings due to CoD [30, 29]. Recent studies show that PINNs can mitigate CoD by leveraging their Monte Carlo nature when PDE solutions are smooth [45]. PINNs often approximate value functions using trainable neural networks by minimizing PDE-driven loss functions, including boundary residual [14, 15], PDE residual [1, 51], and supervised data derived from numerical solvers [28, 50]. Notably, recent studies demonstrate that integrating HJB-based PINNs with Proximal Policy Optimization (PPO) leads to improved performance over standard PPO in continuous-time single-agent MuJoCo environments [27]. However, solving CT-MARL problems through the integration of PINNs and RL remains an open and unexplored area of research.
@@ 3.1 Problem Formulation
In this paper, we focus on multi-agent cooperative settings. Following the continuous-time control system framework [8, 48, 23], we formulate the continuous-time multi-agent problem as a tuple:
@@ 3.2 HJB and Policy Learning
First, we present the following Lemma to show that the optimal value function $V(x,t)$, as defined in Eq. (2), is the optimal solution to HJB PDEs [7] in the context of cooperative multi-agent settings.
@@
To approximate differential value functions, we aim to solve the HJ PDEs given in Eq. (3). However, solving such PDEs through numerical methods becomes computationally intractable for problems with state dimensions higher than 6 [6]. To address this challenge, we adopt PINNs, which approximate value functions to satisfy both PDE residuals and boundary conditions. Specifically, we define HJB PDE residual as
@@ 3.2.2 Policy Learning
Assume the one‑step Q‑function over a short interval $\delta t>0$ be
@@ 3.3 Value Gradient Iteration Module
The performance of continuous-time control policies depends critically on the accuracy of the value, which in turn depends not only on the precision of its own approximation but also on the correctness of the value gradient $\nabla_{x}V(x,t)$. Recent studies have demonstrated that the accuracy of the value directly affects the computational closed-form control inputs [51, 27, 10]. However, the gradients obtained from standard PINN training are often noisy or misaligned with true trajectory behavior. Although prior work [51] improves value gradient approximations using supervised data, such ground-truth gradients are often inaccessible or prohibitively expensive to compute in complex, high-dimensional multi-agent settings. To address this, we introduce the VGI module that iteratively refines the value gradient approximations during training.
@@ 3.4.1 Dynamics Model and Reward Model
In a continuous‑time setting, the true dynamics are given by $\dot{x}=f(x,u)$, but directly learning $f$ via $\frac{x_{t+\Delta t}-x_{t}}{\Delta t}$ as a supervision target is pretty unstable in practice. Instead, we adopt a discrete‑time model‑based approach [36, 20, 13] that we train a neural network $f_{\psi}(x_{t},u_{t},\Delta t)$ to predict the next state $x_{t+\Delta t}$ via
@@
After learning $f_{\psi}$, we recover the continuous‑time derivative by finite differences $\frac{f_{\psi}(x_{t},u_{t},\Delta t)-x_{t}}{\Delta t}$.
@@ 3.4.2 Anchor Loss for Critic Network
Here $\lambda_{\text{term}},\lambda_{\text{anchor}},\lambda_{g}$ are tunable weights balancing PDE fidelity, terminal supervision, value‐bootstrap anchoring, and gradient consistency. In practice, we jointly train the reward model, dynamics model, and value network using the data from the current trajectory. The detailed training process is listed in Appendix A.6.
@@ 4 Experimental Results
We evaluate our Value Iteration via PINN (VIP) method on two continuous‑time multi‑agent benchmarks: MPE [25] and multi‑agent MuJoCo [37, 31]. In addition, we design a didactic benchmark, coupled oscillator (see details in Appendix B.1), to analyze value gradient approximations. This case study is easy to follow and enables the numerical computation of true values and their gradients, which provides a clear and interpretable setting to validate the effectiveness of VIP. Our experiments are designed to answer the following four key questions: (1) Overall efficacy: Does the proposed VIP model outperforms existing continuous‑time RL baselines in these environments? (2) VGI ablation: How much does the VGI module contribute to final performance and training stability? (3) PINN design choice: How does activation function choice and loss term weighting in critic network affect the performance of VIP? (4) Time discretization impact: How well do discrete-time and continuous-time methods perform under arbitrary or unfixed time intervals?
@@ 4.1 Baseline Methods
To evaluate our CT‑MARL framework VIP, we compare against three continuous‑time policy iteration baselines and include an ablated variant of our method without VGI: CT‑MBRL (ODE) [48]: A continuous‑time model‑based RL approach that learns system dynamics via Bayesian neural ODEs. This method uses an actor‑critic framework to approximate state‑value functions and derive the continuous-time optimal policies.
@@
Differential Policy Iteration (DPI) [23]: A model‑based method with differential policy iterations that alternates between (i) solving the HJB PDEs to approximate continuous-time value functions and (ii) updating the policy by following the instantaneous gradient of value approximations. Integral Policy Iteration (IPI) [23]: A partially model‑free approach with integral policy iterations that reformulates the value function as a continuous integral, avoiding explicit differentiation during policy improvement. In our experimental settings, we discretize the integral, roll out trajectories to accumulate rewards, and fit a policy to minimize the resultant value functions. Ablation (w/o VGI): An ablated version of VIP without VGI, which isolates the efficacy of value gradient refinement.
@@ 4.2 Results Analysis
Model Performance. We evaluate all RL methods in MPE and MuJoCo environments using five random seeds and report the mean cumulative reward curves in Fig. 2 and 4. The results show that VIP with VGI consistently converges fastest and achieves the highest final return across all tasks, which empirically validates the efficacy of integrating PINNs with RL. As HJB equations are PDEs with a single boundary condition at terminal time, PINN may struggle to backpropagate the correct physics information when relying solely on boundary values, often resulting in poor value approximations [22, 51].
@@
To address this limitation, we incorporate the anchor and VGI loss terms in Eq. (18), which capture the landscape of value and its gradient so that PINNs converge to the true values. Ablation results further confirm the importance of VGI: removing VGI leads to significantly lower cumulative rewards across all experiments. This observation is consistent with the conclusion in the previous studies [50], which further strengthens that accurate value gradient approximation is crucial for effective PINN training and policy improvement.
@@
Choice of PINN Design. We first examine the impact of activation function choice when incorporating PINN-based losses into critic network refinement. Specifically, we train VIP using ReLU and Tanh activations in both MuJoCo and MPE environments and report the accumulated rewards in Fig. 5. The results show that the VIP with Tanh consistently achieves higher accumulated rewards than the one with ReLU across all tasks. This experiment indicates the importance of activation function choice when using PINN-based losses to refine the critic network and has a consistent conclusion with the previous studies [1, 51]. The good performance of Tanh can be attributed to its smoothness and differentiability, which are particularly important when using PINNs to solve PDEs. PINNs typically use fully-connected network architectures and rely on auto-differentiation to compute value gradients for PDEs. The PINN residuals, including PDEs and boundary conditions, are further optimized using gradient descent. Smooth activation functions like Tanh support stable and accurate gradient flow throughout training, enabling more effective value approximations.
@@
We also investigate the impact of weight parameters for PINN loss terms during VIP training. Specifically, we evaluate three configurations in Eq. (18): 1) balanced weights for all loss terms; 2) a large weight for boundary condition while keeping the others balanced; 3) a large weight for HJB residual with balanced weights for the remaining terms. As shown in Fig. 7, the best accumulated reward performance is achieved when all loss terms are properly balanced. Imbalanced weight settings yield stiffness in the training dynamics of PINNs [42], which makes the dominant loss term (with the largest weights) converge faster than the others [43].
@@
In our experiments, such an imbalance causes VIP only to satisfy the PDEs or boundary conditions during training, ultimately leading to poor value approximations. Lastly, we evaluate the robustness of VIP and MADDPG [24] by generating rollouts with varying time intervals in the didactic environment and computing the average return across these rollouts. Fig. 8 illustrates that VIP maintains a nearly constant return across different time intervals, whereas MADDPG’s performance degrades significantly as the interval increases. This result highlights the advantage of VIP in continuous-time multi-agent scenarios.